Find the parametric equations for the line L through P(5,-2,4) that is parallel to (1/2, 2, -2/3)

(i) where does line L intersect the xy- plane?

(ii) sketch the position vector for a and the line L.


1
Expert's answer
2021-12-14T13:43:41-0500

Parametric equations are


"x=5+\dfrac{1}{2}t, y=-2+2t, z=4-\dfrac{2}{3}t"

(i)


"z=0=>4-\dfrac{2}{3}t=0=>t=6"

"x=5+\dfrac{1}{2}(6)=8, y=-2+2(6)=10"

"Point\ C(8, 10, 0)"


ii)

The position vector is "\vec r=\langle\dfrac{1}{2}, 2, -\dfrac{2}{3}\rangle"